OpenStudy (anonymous):

Convert the equation to the standard form for a parabola by completing the square on x or y as appropriate. y2 + 2y - 2x - 3 = 0 A. (y + 1)2 = 2(x + 2) B. (y - 1)2 = -2(x + 2) C. (y + 1)2 = 2(x - 2) D. (y - 1)2 = 2(x + 2)

OpenStudy (anonymous):

y^2+2 Y-2 X-3=0 Y^2+2 Y+1=2 X+3+1 (Y+1)^2=2 X+4=2(x+2) Which is of the formY^2=4 a X Which is a right formula.

OpenStudy (anonymous):

Oooooh, got it. I feel stupid, I was getting a three at the end for some dumb reason. Thanks!